The highest temperature of this month is (12x-7)°c while the lowest temperature of this month is 2(3x-6)°c. If the difference between the highest and lowest temperature of this month is 23°c, find the highest temperature of the month.

Answer by ikleyn(52781)

.

The governing equation is


    (12x - 7) - 2*(3x-6) = 23


Solve it and find x


    12x - 7 - 6x + 12 = 23

    6x + 5 = 23

    6x     = 23 - 5 = 18

     x              = 18/3

The highest temperature is  (12*3-7) = 36 - 7 = 29 °C.     ANSWER

Solved.